If Jay is 12 years old and Tom is 4 years old. How many years will it be before Tom is 1/2 of Jay's age?
Answer by stanbon(75887)

If Jay is 12 years old and Tom is 4 years old. How many years will it be before Tom is 1/2 of Jay's age?
---
J = 3T
---
Solve: (4+x) = (1/2)(12+x)
8+2x = 12+x
x = 4
---
In 4 yrs Tom will be 8
In 4 yrs Jay will be 16
============================
Cheers,
Stan H.